OpenStudy (anonymous):

2^n+2n is divisible by 4

OpenStudy (goformit100):

yes, do you know the basics of exponents ?

OpenStudy (anonymous):

n belongs to what? N or R

OpenStudy (anonymous):

n= 1 ,2,3 and 4

OpenStudy (anonymous):

no. If n=3 then 2^3+2*3=14.14/4=7/2

OpenStudy (anonymous):

so n= 3 ?

OpenStudy (anonymous):

n= 3 is counterexam for the answer : NO

OpenStudy (anonymous):

the answers i can put are n=1 n=2 n=3 n=4 ... so its not n=3 ?

OpenStudy (anonymous):

It doesn't mean that. It means 2^n+2n is not 'perfectly' divisible by 4 .because if we consider n=3 then there will be a remainder of 2

OpenStudy (anonymous):

so whats n = than

OpenStudy (anonymous):

n={0,1,2,4,........................}. I can not calculate to the infinity.

OpenStudy (anonymous):

i can only pick one i cant pick more then one

OpenStudy (anonymous):

Can you tell me the total question ?

OpenStudy (anonymous):

which of the following is a counterexample for the given statement

OpenStudy (anonymous):

which are the followings...?

OpenStudy (anonymous):

n=1 n=2 n=3 n=4

OpenStudy (anonymous):

which of the following is a counterexample for the given statement n^3+n+1 is a prime

OpenStudy (anonymous):

Is it a new one....?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

n=4

OpenStudy (anonymous):

thanks ....2^n+1 is divisible by 3

OpenStudy (anonymous):

its a new one to

OpenStudy (anonymous):

n=2